a 4.5 kg dog stands on an 18 kg flatboat at distance D ! 6.1 m from the shore. It walks 2.4 m along the boat towardshore and then stops. Assuming no friction between the boat and the wa-ter, find how far the dog is then from the shore.


1
Expert's answer
2021-12-31T08:54:55-0500
"0=m_bv_b-m_dv_d,\\\space\\
v=\dfrac dt,\\\space\\
0=m_b\dfrac{d_b}{t}-m_d\dfrac{d_d}{t},\\\space\\
0=m_bd_b-m_dd_d.\\\space\\
d_b+d_d=2.4 \text{ m.}\\\space\\
d_d=1.92\text{ m},\\
d_d=0.48\text{ m}.\\\space\\
D_\text{dog}=6.1-1.92=4.2\text{ m}."